# Improving Retrieval-Augmented Generation of Large Language Models through Unsupervised General-Purpose Training

### v1

@@
Improving Retrieval-Augmented Generation of Large Language Models through Unsupervised General-Purpose Training
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Unsupervised Training of Large Language Models for Improved Retrieval-Augmented Generation
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Unsupervised Retrieval-Augmented Training of Large Language Models
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Unsupervised Training of Large Language Models as Information Refiner for Retrieval-Augmented Generation
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
InFO-RAG : Information Refinement Training of LLMs for Improved Retrieval-Augmented Generation
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Learning When to Trust the Prefixes: Unsupervised Training for Retrieval-Augmented Generation with
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Learning When to Trust the Prefixes: Unsupervised Information Refinement Training of LLMs for Retrieval-Augmented Generation
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Learning How to Utilize the Input: Unsupervised Information Refinement Training of LLMs for Retrieval-Augmented Generation
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
@@
[…]
${}^{1}$ Key Laboratory of AI Security, Institute of Computing Technology, CAS
${}^{2}$attern Recognition Center, WeChat AI
xushicheng21s@ict.ac.cnpangliang@ict.ac.cnmoyumyu@global.tencent.com
  Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Corresponding authors.
Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) by incorporating additional information from retrieval. However, studies have shown that LLMs still face challenges in effectively using the retrieved information, even ignoring it or being misled by it. The key reason is that the training of LLMs does not clearly make LLMs learn how to utilize input retrieved texts with varied quality. In this paper, we propose a novel perspective that considers the role of LLMs in RAG as “Information Refiner”, which means that regardless of correctness, completeness, or usefulness of retrieved texts, LLMs can consistently integrate knowledge within the retrieved texts and model parameters to generate the texts that are more concise, accurate, and complete than the retrieved texts. To this end, we propose an information refinement training method named InFO [^1] Recursive acronym for InFO reFinement Objective.-RAG that optimizes LLMs for RAG in an unsupervised manner. InFO-RAG is low-cost and general across various tasks. Extensive experiments on zero-shot prediction of 11 datasets in diverse tasks including Question Answering, Slot-Filling, Language Modeling, Dialogue, and Code Generation show that InFO-RAG improves the performance of LLaMA2 by an average of 9.39% relative points. InFO-RAG also shows advantages in in-context learning and robustness of RAG.
[…]
@@ 1 Introduction
(3) Extensive experiments show InFO-RAG enhances the zero-shot RAG of LLaMA2 across Question Answering, Slot-Filling, Language Modeling, Dialog, and Code Generation. InFO-RAG also shows advantages in in-context learning and robustness of RAG. Code will be released at https://github.com/xsc1234/INFO-RAG/.
@@ 3.1 Positive Information Gain in RAG
Scenario 3. The third scenario is that the retrieved texts do not have any answer that can used to solve the question. This scenario means that the question is very difficult or the knowledge involved is very long-tail for information retrieval systems. Even in this case, the retrieval model’s ability to model semantics allows it to provide texts that are semantically related to the question Karpukhin et al. (2020). Therefore, the positive information gain in this scenario is that LLMs can stimulate the knowledge within their parameters based on semantically relevant context to solve the question.
@@ 3.2.2 Data Construction and Training Tasks
In Select and Copy, $s_{l}$ (both $s^{p}_{l}$ and $s^{t}_{l}$) has been contained in the retrieved texts $S$, this needs LLMs to select the texts matching the prefix $s^{p}_{l}$ from the complex retrieved texts $S$ and directly copy the target $s^{t}_{l}$ for generation. The information gain between $s^{t}_{l}$ and input retrieved texts $S$ is that $s^{t}_{l}$ is more concise to be used as the postfix of $s^{p}_{l}$.
For Scenario 2 that needs LLMs to verify the knowledge in the retrieved texts, utilize accurate knowledge, rectify incorrect knowledge, and complete missing knowledge. We propose the training task named Correct and Complete. Given a sentence set $S$, firstly, this task uses the stability of word distribution between layers to get informative tokens. The intention for this is that the more unstable the word distribution of the token among the topmost layers is, the more it indicates that the token is an informative token Chuang et al. (2023). Specifically, for each sentence $s_{i}$ in $S$, our method obtains the next word distribution of the $a\mbox{-}th$ token $s^{[a]}_{i}$ given prefix $s^{<a}_{i}$ of $s_{i}$ in each layer of LLM as:
@@
In Correct and Complete, broken $s_{l}$ with noise is in $S^{\prime}$, LLMs need to extract, correct, and complete the knowledge in $s_{l}$ from $S^{\prime}$ to generate $s^{t}_{l}$.
@@ 4.2 Experimental Settings
As for the retrieval model and retrieval database, for Open-domain QA, Soft Filling and Language Modeling, we use ColBERTv2 Santhanam et al. (2022), a late-interaction model with excellent generalization ability as the retriever, and use Wikipedia consisting of 21,015,324 passages Karpukhin et al. (2020) as retrieval database. For Code Generation, we SCODE-R Parvez et al. (2021) as code retriever and use deduplicated source codes in CodeSearchNET Husain et al. (2019) as retrievel database. For all the above tasks, we give Top-5 retrieved passages to each example. For LFQA, Dialog, and Multi-Hop QA, we use the list of contextual passages provided in the datasets as the retrieved list (distractor setting).
@@ A.3 Robustness to Number of Retrieved Passages
A.4 Performance on MMLU
@@
Humanities STEM Social Sciences Other Average

### v2

# Unsupervised Information Refinement Training of Large Language Models for Retrieval-Augmented Generation

@@
Abstract
Retrieval-augmented generation (RAG) enhances large language models (LLMs) by incorporating additional information from retrieval. However, studies have shown that LLMs still face challenges in effectively using the retrieved information, even ignoring it or being misled by it. The key reason is that the training of LLMs does not clearly make LLMs learn how to utilize input retrieved texts with varied quality. In this paper, we propose a novel perspective that considers the role of LLMs in RAG as “Information Refiner”, which means that regardless of correctness, completeness, or usefulness of retrieved texts, LLMs can consistently integrate knowledge within the retrieved texts and model parameters to generate the texts that are more concise, accurate, and complete than the retrieved texts. To this end, we propose an information refinement training method named InFO-RAG that optimizes LLMs for RAG in an unsupervised manner. InFO-RAG is low-cost and general across various tasks. Extensive experiments on zero-shot prediction of 11 datasets in diverse tasks including Question Answering, Slot-Filling, Language Modeling, Dialogue, and Code Generation show that InFO-RAG improves the performance of LLaMA2 by an average of 9.39% relative points. InFO-RAG also shows advantages in in-context learning and robustness of RAG.
@@ Abstract
Shicheng Xu1,2thanks:   Work done during the Tencent Rhino-bird Research Elite Program at WeChat.  Liang Pang1thanks:   Corresponding authors.  Mo Yu322footnotemark: 2  Fandong Meng3  Huawei Shen1  Xueqi Cheng1  Jie Zhou3 1CAS Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2University of Chinese Academy of Sciences 3Pattern Recognition Center, WeChat AI {xushicheng21s,pangliang,shenhuawei,cxq}@ict.ac.cn moyumyu@global.tencent.com  {fandongmeng,withtomzhou}@tencent.com
@@ 1 Introduction
(3) Extensive experiments show InFO-RAG enhances the zero-shot RAG of LLaMA2 across Question Answering, Slot-Filling, Language Modeling, Dialog, and Code Generation. InFO-RAG also shows advantages in in-context learning and robustness of RAG. Code is released at https://github.com/xsc1234/INFO-RAG/.
@@ 3.1 Positive Information Gain in RAG
Scenario 3. The third scenario is that the retrieved texts do not have any answer that can used to solve the question. This scenario means that the question is very difficult or the target knowledge is very long-tail for information retrieval systems. Even in this case, the retrieval model’s ability to model semantics allows it to provide texts that are semantically related to the question Karpukhin et al. (2020). Therefore, the positive information gain in this scenario is that LLMs can stimulate the knowledge within their parameters based on semantically relevant context to solve the question.
@@ 3.2.2 Data Construction and Training Tasks
In Select and Copy, $s_{l}$ (both $s^{p}_{l}$ and $s^{t}_{l}$) has been contained in the retrieved texts $S$, this needs LLMs to select the texts matching the prefix $s^{p}_{l}$ from the complex retrieved texts $S$ and directly copy the target $s^{t}_{l}$ for generation. The information gain between $s^{t}_{l}$ and input retrieved texts $S$ is that $s^{t}_{l}$ is more concise to be used as the postfix for $s^{p}_{l}$.
For Scenario 2 that needs LLMs to verify the knowledge in the retrieved texts, utilize accurate knowledge, rectify incorrect knowledge, and complete missing knowledge. We propose the training task named Correct and Complete. Given a sentence set $S$, firstly, this task uses the stability of word distribution between layers to get informative tokens. The intention for this is that the more unstable the word distribution of the token among the topmost layers is, the more it indicates that the token is an informative token. We follow Chuang et al. (2023) to achieve this. Specifically, for each sentence $s_{i}$ in $S$, our method obtains the next word distribution of the $a\mbox{-}th$ token $s^{[a]}_{i}$ given prefix $s^{<a}_{i}$ of $s_{i}$ in each layer of LLM as:
@@
In Correct and Complete, the broken $s_{l}$ with noise is already in $S^{\prime}$. The information gain in this task requires LLM to extract, correct, and complete the knowledge in $s_{l}$ from $S^{\prime}$ to generate $s^{t}_{l}$.
@@ 4.2 Experimental Settings
As for the retrieval model and retrieval database, for Open-domain QA, Soft Filling and Language Modeling, we use ColBERTv2 Santhanam et al. (2022), a late-interaction model with excellent generalization ability as the retriever, and use Wikipedia consisting of 21,015,324 passages Karpukhin et al. (2020) as retrieval database. For Code Generation, we SCODE-R Parvez et al. (2021) as code retriever and use deduplicated source codes in CodeSearchNET Husain et al. (2019) as retrievel database. For all the above tasks, we give Top-5 retrieved passages to each example. For LFQA, Dialog, and Multi-Hop QA, we use the list of contextual passages provided in the datasets as the retrieved list (distractor setting). In each experiment, all baselines and our method share the same retrieved documents.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Key R&D Program of China (2022YFB3103700, 2022YFB3103704), the National Natural Science Foundation of China (NSFC) under Grants No. 62276248 and U21B2046, and the Youth Innovation Promotion Association CAS under Grants No. 2023111.
@@ A.3 Robustness to Number of Retrieved Passages
T-REx ZS NQ WebQ
Baseline 51.47 40.26 45.05 41.78
+ INFO-RAG 55.67 43.29 49.76 44.02
Table 11: Works based on BM25.
A.4 Ablation Study on Masking Strategy
T-REx ZS NQ WebQ
Baseline 75.96 43.79 5.59 67.03 16.58 1.42 69.37 30.72 6.16 65.07 31.88 5.47
Simple Mask 78.43 44.05 5.75 70.30 19.45 1.96 73.59 31.05 6.51 70.55 32.96 6.83
Our method 79.25 48.59 6.67 70.26 25.02 3.87 73.73 33.85 8.39 70.59 37.48 11.25
Table 12: Ablation study of masking strategy on three scenarios. “has-ans.” is the first scenario that correct answers are in retrieved texts. “replace” is the second scenario that correct answers are randomly replaced with other phrases to simulate the incorrect and incomplete knowledge. “no-ans.” is the third scenario that retrieval cannot find any answers.
In general, Table 13 and 12 show our masking strategy in Scenario 3 is more effective than simple and straightforward masking. Specifically, our method is more significantly effective in the scenarios that correct answers are randomly replaced with other phrases (replace) and retrieval cannot find any answers (no answer).
A.5 Works with Different Retriever
We evaluate our method and baseline (LLaMA2-13B-chat) with BM25 as the retriever, the experimental results shown in Table 11 indicate that our method still performs better than baseline when the retriever as BM25.
Method T-REx ZS NQ WebQ
Baseline 62.53 56.81 50.36 45.47
Simple Mask 64.05 58.91 53.80 50.55
Our method 65.39 59.05 54.04 51.07
Table 13: Ablation study of masking strategy.
A.6 Performance on MMLU
@@
Humanities STEM Social-Sciences Other Average